what is the quadratic equation of 8y+3y^2=35? please help!
Answer by waynest(281)

8y + 3y^2 = 35
3y^2 + 8y - 35 = 0
use quadratic equation to solve:
(-8 +-sqrt8^2 - 4(3)(-35))/2(3)
(-8 +-sqrt64 +420) /6
(-8 +-sqrt484) /6
(-8 +- 22) /6
(-4 +- 11) /3
x = (-4 + 11) /3 = 2.33
x = (-4 - 11) /3 = -5